Suppose g : A → B and f : B → C are functions.
a. Show that if f ◦g is onto, then f must also be onto.
b. Show that if f ◦g is one-to-one, then g must also be one-to-one.
c. Show that if f ◦g is a bijection, then g is onto if and only if f is one-to-one.
1
Expert's answer
2019-09-19T12:24:40-0400

Suppose "g : A\to B" and "f : B\to C"  are functions.

a. Show that if "f\circ g"  is onto, then "f"  must also be onto.

Answer. Assume that "f\circ g"  is onto. Let "c\in C" . Since "f\circ g"  is onto, there is "a\in A"  such that "c = (f\circ g)(a)"  by the definition of onto function. By the definition of function composition, "c = f(g(a))" . Let "b = g(a)" . This "b"  belongs to "B"  and is such that "c = f(b)". Therefore, for every "c\in C" , there is "b\in B"  such that "c = f(b)" . By the definition of onto function, "f"  is onto.

b. Show that if "f\circ g"  is one-to-one, then "g"  must also be one-to-one.

Answer. Let "a_1, a_2\in A"  such that "g(a_1) = g(a_2)" . Then

"(f\circ g)(a_1) = f(g(a_1)) = f(g(a_2))= (f\circ g)(a_2)."

 Since "f\circ g"  is one-to-one, "a_1=a_2" . Therefore, for all "a_1, a_2\in A"  such that "g(a_1) = g(a_2)" , "a_1=a_2" . By the definition of one-to-one function, "g"  is one-to-one.

c. Show that if "f\circ g"  is a bijection, then "g"  is onto if and only if "f"  is one-to-one.

Answer. Assume that "f\circ g"  is a bijection.

Assume that "g"  is onto. Let "b_1, b_2\in B"  such that "f(b_1) = f(b_2)" . Since "g"  is onto, by the definition of onto function, there are "a_1, a_2\in A"  such that "b_1 = g(a_1)"  and "b_2 = g(a_2)" . We have

"(f\circ g)(a_1) = f(g(a_1)) = f(b_1)"

"= f(b_2) = f(g(a_2)) = (f\circ g)(a_2)."

 Since "f\circ g"  is a bijection, it is one-to-one, hence "a_1=a_2" , and

"b_1 = g(a_1) = g(a_2) = b_2."
  •  Therefore, for all "b_1, b_2\in B"  such that "f(b_1) = f(b_2)" , "b_1=b_2" . By the definition of one-to-one function, "f"  is one-to-one.
  • Assume that "f"  is one-to-one. Let "b\in B" . Since "f\circ g"  is a bijection, it is onto, hence there is "a\in A"  such that "f(b) = (f\circ g)(a) = f(g(a))" . Since "f"  is one-to-one, "b = g(a)" . Therefore, for every "b\in B" , there is "a\in A"  such that "b = g(a)" . By the definition of onto function, "g"  is onto.